# The microscale organization of directed hypergraphs

### v1

@@ Abstract
Many real-world complex systems are characterized by non-pairwise – higher-order – interactions among system’s units, and can be effectively modeled as hypergraphs. Directed hypergraphs distinguish between source and target sets within each hyperedge, and allow to account for the directional flow of information between nodes. Here, we provide a framework to characterize the structural organization of directed higher-order networks at their microscale. First, we extract the fingerprint of a directed hypergraph, capturing the frequency of hyperedges with a certain source and target sizes, and use this information to compute differences in higher-order connectivity patterns among real-world systems. Then, we formulate reciprocity in hypergraphs, including exact, strong, and weak definitions, to measure to which extent hyperedges are reciprocated. Finally, we extend motif analysis to identify recurring interaction patterns and extract the building blocks of directed hypergraphs. We validate our framework on empirical datasets, including Bitcoin transactions, metabolic networks, and citation data, revealing structural principles behind the organization of real-world systems.
@@ Introduction
Accurately modeling interactions among entities is crucial to understand the properties of many complex systems. Traditional network models focus on pairwise connections between nodes [1, 2], neglecting the complexities of systems where multiple units interact simultaneously. Such higher-order interactions are prevalent in various domains, including social networks [3, 4], folksonomies [5], ecological systems [6], chemical reactions [7] including metabolic pathways [8], and the brain [9, 10].
@@
Hypergraphs [11] provide a framework for explicitly encoding higher-order interactions, representing them as hyperedges connecting multiple nodes simultaneously. By preserving group-based interactions, they improve our ability to understand the structures and dynamics of systems with many-body interactions [12, 13]. Recently, a variety of measures have been introduced or extended to capture the higher-order organization of complex systems, including centrality [14, 15], community structure [16, 17, 18] and motifs [19, 20, 21]. Moreover, new models have allowed to describe systems’ evolution [22, 23, 24], and highlight the importance of higher-order interactions in shaping emergent behaviors in diffusion [25, 26], synchronization [27, 28, 29], spreading [30, 31] and evolutionary dynamics [32].
@@
Most research has so far focused on undirected hypergraphs, which fail to capture the directional nature of many real-world interactions. For example, in a metabolic reaction, a set of reactants transforms into a set of products [8]. Similarly, in a Bitcoin transaction, multiple source wallets may transfer funds simultaneously to multiple target wallets [33]. To accurately encode such interactions, models must incorporate directionality into their representations. In this sense, directed hypergraphs enhance modeling by distinguishing between source and target sets in each hyperedge [34]. Tools to study directed hypergraphs are largely underdeveloped, with notable exceptions in areas such as null models [35], synchronization [36], overlapping patterns between two hyperedges of limited size [37], and some early proposals to define reciprocity [38, 39].
@@
In this work, we introduce measures and tools to characterize the microscale organization of real-world directed hypergraphs. First, we discuss a decomposition into fundamental interaction types: one-to-one, one-to-many, many-to-one and many-to-many. We analyze empirical data to count the occurrences of each interaction type, and use this information as a signature to compute differences in higher-order connectivity patterns. Second, we propose new simple and computationally efficient definitions for reciprocity [40] for directed hypergraphs, namely exact, strong and weak higher-order reciprocity, designed to capture different patterns of bi-directionality in empirical data. Finally, we extend motif analysis [41] to incorporate the directionality of interactions, extracting recurring higher-order and directed subgraphs. Our results suggest the existence of complex mechanisms of feedback and reinforcement in the information flow among system units, where pairwise interactions support the action of groups, and vice versa.
@@ Results
Traditional graph models reduce directed group interactions into a collection of pairwise links, often leading to a loss of important structural information about group organization and dynamics. For instance, reducing a many-to-many interaction such as SOURCE = $\{A,B\}$ and TARGET = $\{D,E\}$ to a set of pairwise directed links ($A\to D$, $A\to E$, $B\to D$ and $B\to D$) fails to capture the collective nature of the interaction, where pairs of nodes are jointly involved in the source and target sets. Directed hypergraphs preserve group-based structure, allowing for a more faithful representation of complex interactions. In such a mathematical framework, hyperedge direction is encoded by distinguishing between source and target node sets, which are non-empty and disjoint. In particular, we distinguish four fundamental patterns of interactions encoded as directed hyperedges: one-to-one, where a single source node connects to a single target; one-to-many, where one source affects multiple targets; many-to-one, where multiple sources act on a single target; and many-to-many, the most general case, where multiple sources act on multiple targets. In Figure 1, we show an example of a directed hypergraph, highlighting all distinct hyperedge patterns.
To study the microscale organization of real-world systems with directed group interactions, we collected datasets from multiple domains and mapped them into directed hypergraphs. The datasets [39] include qna (nodes are users and forum posts are hyperedges), E-mail (nodes are users and emails are hyperedges), Bitcoin (nodes are accounts and financial transactions are hyperedges), Metabolic (nodes are genes and metabolic reactions are hyperedges) and Citation (nodes are authors and hyperedges are paper citations). Detailed descriptions and summary statistics of each dataset are reported in Supplementary Note 1.
@@
Figure 2: Hyperedge signature of directed hypergraphs. a) We characterize each system with a signature hyperedge vector, encoding the abundance of a certain pattern of directed hyperedge. Systems within the same domain share the same color. b) Dendrogram resulting from agglomerative clustering applied to the correlation matrix of hyperedge signature vectors for each dataset. Correlation values are color-coded, with high positive correlations in red and high negative correlations in blue.
@@
We characterize directed hypergraphs across domains by investigating the diversity in their patterns of directed hyperedges. For each dataset, we construct a hyperedge signature vector $\mathbf{v}$, which captures the distribution of hyperedges based on the sizes of their source and target sets (see Methods). Such vectors provide a fingerprint for systems based on their higher-order connectivity patterns at the microscale. Figure 2a shows the hyperedge signature vectors for each dataset, considering interactions up to size $6$. To emphasize the role of higher-order interactions in the analysis, we do not consider one-to-one interactions. In the E-mail data we find abundance only in entries corresponding to one-to-many interactions, reflecting the typical structure of email communications. Similarly, in the qna, many-to-one interactions are prevalent, as these systems involve multiple individuals responding to a question by a single user. In contrast, Metabolic and Citation datasets show high abundances in many-to-many relationships across a variety of source and target set sizes. Finally, Bitcoin dataset exhibits more varied behavior, with abundant entries for both one-to-many and many-to-many interactions, indicating different interaction types in the network.
Figure 3: Reciprocity measures for directed hypergraphs. For exact reciprocity, the direction of a hyperedge is fully reversed by a single hyperedge in which the source and target sets are swapped. In strong reciprocity, multiple hyperedges collectively reverse the interaction, with the source and target sets being fully reciprocated through a combination of interactions. In weak reciprocity, at least one node from the target set reciprocates an interaction with one node from the source set.
@@
To further explore structural diversity across different domains, we compute pairwise correlations between hyperedge signature vectors (Pearson coefficient) and apply hierarchical agglomerative clustering on their correlation matrix. A correlation value close to $1$ indicates similar hyperedge structures, 00 suggests no relationship, and $-1$ indicates the structures are inversely related. The clustering procedure applied to the systems’ correlation matrix results in a dendrogram that visually represents their hierarchical relationships, highlighting the presence of clusters of directed hypergraphs that share similar connectivity patterns. In Figure 2b, we show the correlation matrix and the clustering dendrogram. By examining the correlation matrix, we observe a strong correlation within systems from the same domain, indicating highly similar abundance in hyperedge structures. In contrast, systems from different domains exhibit varying degrees of correlation. Specifically, E-mail and qna datasets are inversely correlated, as they display non-overlapping and complementary connectivity patterns: E-mail is characterized by one-to-many interactions, whereas qna primarily involves many-to-one relationships. The Metabolic and Citation datasets, which feature many-to-many interactions, are positively correlated and form a distinct cluster. Interestingly, the Bitcoin datasets also display positive correlations with the Metabolic and Citation cluster due to a high presence of many-to-many interaction patterns. However, they also exhibit a weaker positive correlation with the E-mail datasets, reflecting the presence of one-to-many interactions in Bitcoin.
@@
Reciprocity is a fundamental property of systems with directed interactions, including social networks [42]. It traditionally refers to the tendency of the system’s units to mutually exchange directed interactions. In directed graphs, reciprocity is defined as $r=\frac{\overleftrightarrow{L}}{L}$, i.e., the ratio of the number of bidirectional links ($\overleftrightarrow{L}$) to the total number of links ($L$). This measure has been widely used to describe real-world directed networks [43, 40]. Recognizing its broad importance, recent works have extended reciprocity to hypergraphs, accounting for the complexity of having multiple nodes in both the source and target sets of hyperedges. Among the recent approaches for hypergraph reciprocity, one method decomposes hyperedges into pairwise links [38], losing information about group interactions. An alternative approach defines a more complex measure that diverges from the traditional binary definition of reciprocity at the level of single links [39]. While this can capture finer nuances, such a measure is computationally expensive and more difficult to interpret.
Figure 4: Higher-order reciprocity in real-world hypergraphs. Number of reciprocated hyperedges for each different notion of reciprocity. In blue total hyperedges, in yellow exactly reciprocated hyperedges, in green strongly reciprocated hyperedges and in orange weakly reciprocated hyperedges. Statistics are disaggregated by hyperedge size. To simplify the plot, we grouped higher-order reciprocity of systems from the same domain.
@@
  • Exact reciprocity occurs when an interaction represented by a hyperedge with a source set $h$ and a target set $t$ is precisely mirrored by another interaction with the source and target sets reversed. Formally, two hyperedges $e_{1}=(h_{1},t_{1})$ and $e_{2}=(h_{2},t_{2})$ are exactly reciprocated if and only if $h_{1}=t_{2}\text{ and }t_{1}=h_{2}$. This is the strictest form of reciprocity.
@@
  • Weak reciprocity represents the most relaxed form of reciprocity and requires only that at least one node from the target set of a hyperedge appears in the source set of another, and vice versa. Formally, a hyperedge $e=(h,t)$ is weakly reciprocated if there exists another hyperedge $e^{\prime}=(h^{\prime},t^{\prime})$ such that $h\cap t^{\prime}\neq\emptyset\text{ and }t\cap h^{\prime}\neq\emptyset$.
We summarize our definitions of reciprocity for directed hypergraphs in Figure 3. In Figure 4, we show statistics about the number of reciprocated hyperedges in each domain and for different definitions of reciprocity. Hyperedges are further disaggregated by size. Overall, every system displays a certain degree of reciprocity. In particular, weak reciprocity is widespread across domains, and the proportion of weakly reciprocated hyperedges remains relatively constant regardless of hyperedge size. Strongly reciprocated hyperedges are also common in each system, but their proportion decreases with hyperedge size at different domain-dependant rates. Conversely, exact reciprocity is rare. In particular, qna and E-mail prevent exactly reciprocated hyperedges (e.g., a single e-mail cannot have more senders simultaneously). Instead, Citation hypergraphs exhibit the highest proportion of exactly reciprocated hyperedges, observed across all hyperedge sizes. Notably, Bitcoin and qna display the lowest levels of reciprocity across all definitions, and their degree of reciprocity is more sensitive to hyperedge size than in other systems.
@@
Motif analysis involves counting the frequency of patterns of interactions in connected subgraphs of a given number of nodes. This framework was first introduced by Milo et al. [41] to extract the fundamental functional units of complex systems [44]. Recently, motif analysis has been extended to hypergraphs to capture patterns of interactions with arbitrary size [19]. Here, we extend such analysis to consider also the direction of the hyperedges involved in the patterns.
@@
First, it is interesting to study the combinatorics of the patterns of directed subhypergraphs. There is no simple closed-form formula for counting the number of possible directed higher-order motifs as a function of their order $n$, i.e., the number of nodes in the patterns. We can estimate the number of non-isomorphic connected directed hypergraphs in a way similar to [19]. Given a set of $n$ nodes, the number of possible directed hyperedges is $13^{n}-2\cdot\sum_{k=1}^{n}\binom{n}{k}-1=3^{n}-2\cdot 2^{n}+$ 1. This expression counts the ways to partition the $n$ nodes into three disjoint sets: source, target and empty set. We subtract the invalid combinations with empty source or target sets. Given $n$ nodes, we ensure connectivity by selecting a chain of $n-1$ hyperedges and including them in the hypergraph, leaving us with $23^{n}-2\cdot 2^{n}-n+$ 2 remaining possible hyperedges. For each remaining hyperedge, we decide whether to include it or not, resulting in $2^{3^{n}-2\cdot 2^{n}-n+2}$ total hypergraphs. Since we are interested in non-isomorphic hypergraphs, we divide this number by $n!$, the number of ways to label the vertices, providing the lower bound $\frac{2^{3^{n}-2\cdot 2^{n}-n+2}}{n!}$. If we ignore the constraints of non-isomorphism and connectivity, we count the number of possible labeled hypergraphs. Since each of the $13^{n}-2\cdot 2^{n}+$ 1 possible hyperedges can either be included or excluded, the total number of labeled hypergraphs is at most $2^{3^{n}-2\cdot 2^{n}+1}$. Figure 5 shows the upper and lower bounds on the growth of possible sub-hypergraph patterns as a function of the number of nodes (order), for both the undirected and directed cases. The estimated number of patterns grows super-exponentially, even in the undirected case. In the directed case, the growth is even faster due to the need to consider all possible subdivisions into source and target sets.

### v2

@@ Abstract
Many real-world complex systems are characterized by non-pairwise – higher-order – interactions among system’s units, and can be effectively modeled as hypergraphs. Directed hypergraphs distinguish between source and target sets within each hyperedge, and allow to account for the directional flow of information between nodes. Here, we provide a framework to characterize the structural organization of directed higher-order networks at their microscale. First, we extract the fingerprint of a directed hypergraph, capturing the frequency of hyperedges with a certain source and target sizes, and use this information to compute differences in higher-order connectivity patterns among real-world systems. Then, we investigate the overlap among sources and targets to reveal recurring sets of co-sending and co-receiving nodes. We define reciprocity in hypergraphs using exact, strong, and weak definitions to quantify the extent to which hyperedges are reciprocated. Finally, we extend motif analysis to identify recurring interaction patterns and extract the building blocks of directed hypergraphs. We validate our framework on empirical datasets, including Bitcoin transactions, metabolic networks, and citation data, revealing structural principles behind the organization of real-world systems.
@@ Introduction
Accurately modeling interactions among entities is crucial to understand the properties of many complex systems. Traditional network models focus on pairwise connections between nodes [boccaletti2006complex, cimini2019statistical], neglecting the complexities of systems where multiple units interact simultaneously. Such higher-order interactions are prevalent in various domains, including social networks [patania2017shape, cencetti2021temporal, iacopini2024temporal], folksonomies [ghoshal2009random], ecological systems [grilli2017higher], chemical reactions [jost2019hypergraph] including metabolic pathways [traversa2023robustness], and the brain [petri2014homological, santoro2023higher].
@@
Hypergraphs [berge1973graphs] provide a framework for explicitly encoding higher-order interactions, representing them as hyperedges connecting multiple nodes simultaneously. By preserving group-based interactions, they improve our ability to understand the structures and dynamics of systems with many-body interactions [battiston2020networks, battiston2021physics]. Recently, a variety of measures have been introduced or extended to capture the higher-order organization of complex systems, including centrality [benson2019three, tudisco2021node], community structure [eriksson2021choosing, contisciani2022inference, ruggeri2023community] and motifs [lotito2022higher, lee2020hypergraph, arregui2024patterns]. Moreover, new models have allowed to describe systems’ evolution [petri2018simplicial, digaetano2024percolation, gallo2024higher], and highlight the importance of higher-order interactions in shaping emergent behaviors in diffusion [schaub2020random, carletti2020random], synchronization [lucas2020multiorder, gambuzza2021stability, zhang2023higher], spreading [iacopini2019simplicial, chowdhary2021simplicial] and evolutionary dynamics [civilini2024explosive].
@@
Most research has so far focused on undirected hypergraphs, which fail to capture the directional nature of many real-world interactions. For example, in a metabolic reaction, a set of reactants transforms into a set of products [traversa2023robustness]. Similarly, in a Bitcoin transaction, multiple source wallets may transfer funds simultaneously to multiple target wallets [ranshous2017exchange]. To accurately encode such interactions, models must incorporate directionality into their representations. In this sense, directed hypergraphs enhance modeling by distinguishing between source and target sets in each hyperedge [gallo1993directed]. Tools to study directed hypergraphs are largely underdeveloped, with notable exceptions in areas such as null models [preti2024higher], synchronization [gallo2022synchronization], overlapping patterns between two hyperedges of limited size [moon2023four], and some early proposals to define reciprocity [pearcy2014hypergraph, kim2023reciprocity].
@@
In this work, we introduce measures and tools to characterize the microscale organization of real-world directed hypergraphs. First, we discuss a decomposition into fundamental interaction types: one-to-one, one-to-many, many-to-one and many-to-many. We analyze empirical data to count the occurrences of each interaction type, and use this information as a signature to compute differences in higher-order connectivity patterns. Then, for each node, we investigate the overlap among its source and target sets, to extract recurring groups of co-senders and co-receivers. By examining this overlap and comparing it against randomised models, we aim to reveal whether certain systems exhibit a more redundant organization, where interactions frequently recur among the same groups, or a more diverse structure with less overlap among participants. Additionally, we propose new, computationally efficient definitions for reciprocity [garlaschelli2004patterns] for directed hypergraphs, namely exact, strong and weak higher-order reciprocity, designed to capture different patterns of bi-directionality in empirical data. Finally, we extend motif analysis [milo2002network] to incorporate the directionality of interactions, extracting recurring higher-order and directed subgraphs. Our results suggest the existence of complex mechanisms of feedback and reinforcement in the information flow among system units, where pairwise interactions support the action of groups, and vice versa.
@@ Results
[…]
We analyzed datasets from multiple domains, including qna (nodes are users and forum posts are hyperedges), E-mail (nodes are users and emails are hyperedges), Bitcoin (nodes are accounts and financial transactions are hyperedges), Metabolic (nodes are genes and metabolic reactions are hyperedges) and Citation (nodes are authors and hyperedges are paper citations) [kim2023reciprocity]. Each dataset is encoded as a set-indexed adjacency tensor. In particular, we index the distinct source- and target-sets observed in the data by $\{S_{\alpha}\}_{\alpha=1}^{p}$ and $\{T_{\beta}\}_{\beta=1}^{q}$, respectively, and define the set-indexed adjacency tensor $\mathcal{A}\in\{0,1\}^{p\times q}$ by $\mathcal{A}_{\alpha\beta}=1$ if and only if there exists $e\in E$ with $\mathsf{s}(e)=S_{\alpha}$ and $\mathsf{t}(e)=T_{\beta}$ (and $\mathcal{A}_{\alpha\beta}=0$ otherwise). Whenever $\mathcal{A}_{\alpha\beta}=1$, we enforce $S_{\alpha}\neq\varnothing$, $T_{\beta}\neq\varnothing$ and $S_{\alpha}\cap T_{\beta}=\varnothing$.
Detailed descriptions and summary statistics of each dataset are reported in Supplementary Note 1.
@@
Figure 2: Hyperedge signature of directed hypergraphs. a) We describe each system with a hyperedge signature vector whose entries encode the count of directed hyperedges with source and tail sizes $(|S|,|T|)$. We compute statistics using hyperedges with total cardinality at most 66 (i.e., $|S|+|T|\leq 6$). For visualization, we display each vector as a sequence of histogram panels: one panel for each source size $|S|=i$, separated by a small gap; within each panel, bins correspond to tail sizes $|T|$ in increasing order, restricted by $i+|T|\leq 6$. Systems from the same domain share the color. b) Dendrogram resulting from agglomerative clustering applied to the correlation matrix of hyperedge signature vectors for each dataset. Correlation values are color-coded, with high positive correlations in red and high negative correlations in blue.
@@
A natural starting point to characterize directed hypergraphs across domains is investigating the diversity in their patterns of directed hyperedges. For each dataset, we construct a hyperedge signature vector $\mathbf{v}$, which captures the distribution of hyperedges based on the sizes of their source and target sets (see Methods). Such vectors provide a fingerprint for systems based on their higher-order connectivity patterns at the microscale. Figure 2a shows the hyperedge signature vectors for each dataset, considering interactions up to size $6$. To emphasize the role of higher-order interactions in the analysis, we do not consider one-to-one interactions. We find that one-to-many interactions dominate the E-mail dataset, reflecting the typical structure of email communications. Similarly, in the qna, many-to-one interactions are prevalent, as these systems involve multiple individuals responding to a question by a single user. In contrast, Metabolic and Citation datasets show high abundances in many-to-many relationships across a variety of source and target set sizes. Finally, Bitcoin dataset exhibits more varied behavior, with abundant entries for both one-to-many and many-to-many interactions, indicating different interaction types in the network.
@@
To further explore structural diversity across different domains, we compute pairwise rank correlations among hyperedge signature vectors using weighted Kendall’s $\tau$ and apply hierarchical agglomerative clustering on their correlation matrix. A correlation value close to $1$ indicates similar hyperedge structures, 0 suggests no relationship, and 1-1 indicates the structures are inversely related. The clustering procedure applied to the systems’ correlation matrix results in a dendrogram that visually represents their hierarchical relationships, highlighting the presence of clusters of directed hypergraphs that share similar connectivity patterns. In Figure 2b, we show the correlation matrix and the clustering dendrogram. By examining the correlation matrix, we observe a strong correlation within systems from the same domain, indicating highly similar abundance in hyperedge structures. In contrast, systems from different domains exhibit varying degrees of correlation. Specifically, E-mail and qna datasets are inversely correlated, as they display non-overlapping and complementary connectivity patterns: E-mail is characterized by one-to-many interactions, whereas qna primarily involves many-to-one relationships. The Metabolic and Citation datasets, which feature many-to-many interactions, are positively correlated and form a distinct cluster. Interestingly, the Bitcoin datasets also display positive correlations with the Metabolic and Citation cluster due to a high presence of many-to-many interaction patterns. However, they also exhibit a weaker positive correlation with the E-mail datasets, reflecting the presence of one-to-many interactions in Bitcoin.
Source and target sets overlap
Figure 3: Overlap across domains. a) Distribution of node counts within the joint z-score space of source and target overlap, representing how much nodes deviate from null model expectations in both dimensions. b) Bar plots quantifying the fraction of nodes exceeding the $\textit{z-score}\geq 2$ threshold for either source or target overlaps, nodes exceeding both thresholds, and nodes below both thresholds.
The degree to which hyperedges share elements provides valuable insights into redundancy, hierarchical structures, and information flow within the system. In general, real-world systems exhibit a high degree of overlap, indicating that recurrent and redundant interactions are a shared feature [lee2021hyperedges]. Moreover, hyperedge overlap has been shown to widely impact the dynamics of systems with higher-order interactions [malizia2025hyperedge, lamata2025hyperedge].
In directed hypergraphs, nodes are frequently involved in multiple hyperedges, either as part of the source or the target set. In order to characterize and quantify this property, for each node, we measure how much its incident source sets overlap and how much its incident target sets overlap, and we compare with a null model, reporting source and target z-scores (see Methods). In other words, we measure the extent to which nodes engage in interactions with the same set of co-senders or co-receivers. Specifically, an observed overlap associated with a z-score greater than $2$ (i.e., statistically significant excess overlap) highlights nodes that tend to participate in structurally redundant interactions where the same groups of nodes frequently co-occur in source or target sets. Conversely, a z-score less than 2-2 reflects statistically significant lower overlap than expected, suggesting that interactions are more diverse, with hyperedges being more distinct and less likely to share members.
In Fig. 3, we show the distribution of nodes with a given overlap z-score for source and target sets across domains, along with the fraction of nodes in each region of the z-score space. The focus is on positive excess overlap, as statistically significant negative excess overlap is very rare. The Citation dataset displays significant overlap for both source and target sets, showing that (i) an author tends to preferentially work with known collaborators, and (ii) an author tends to be cited repeatedly alongside similar sets of authors. In the E-mail dataset, the excess overlap can be computed only for the target sets, as the source sets always have cardinality $1$. The overlap is significantly larger than random, underscoring the hierarchical and broadcast-like nature of email communication. The Bitcoin dataset generally exhibits nodes with high excess overlap in both source and target sets. However, the presence of nodes with excess overlap values lower than zero implies that certain participants in the network engage in interactions that introduce more novelty rather than reinforce existing hyperedges. The Metabolic dataset follows a similar trend, with half of the nodes displaying significant excess overlap for either source or target sets or both, suggesting that metabolic reactions tend to involve recurring sets of substrates and products and highlighting the modular nature of metabolic networks. qna data show a lower excess overlap compared with the other systems, indicating that forum respondents are less likely to engage repeatedly with the same set of co-responders. Since the target sets in this dataset always have cardinality $1$, the excess overlap can be computed only for source sets.
@@
Figure 4: Reciprocity measures for directed hypergraphs. a) Example of a directed hyperedge. b) All possible ways in which this hyperedge can be reciprocated according to our definitions. Exact reciprocity: a single hyperedge with source and target sets swapped, represented by reversing the arrow between the same node sets. Strong reciprocity: multiple hyperedges collectively reverse the interaction, possibly involving external nodes. Weak reciprocity: at least one node in the target set reciprocates with one node in the source set, illustrated as a pairwise link with reversed direction. In all panels, shaded areas group nodes involved in each interaction; colors encode the interaction pattern (green many-to-many, red many-to-one, black one-to-one); arrows encode direction from source to target. Grey disks denote nodes in the original hyperedge, and external nodes that appear only in reciprocal interactions are white with a dashed border.
Reciprocity is a fundamental property of systems with directed interactions, including social networks [wasserman1994social]. It traditionally refers to the tendency of the system’s units to mutually exchange information. In directed graphs, reciprocity is defined as
$r=\frac{\overleftrightarrow{L}}{L}$
, measuring the ratio of the number of bidirectional links ($\overleftrightarrow{L}$) to the total number of links ($L$). This measure is often normalized as
$\rho_{\mathrm{NM}}=\frac{r-\langle r\rangle_{\mathrm{NM}}}{1-\langle r\rangle_{\mathrm{NM}}}$
where $r$ is the observed reciprocity in the network, and $\langle r\rangle_{\mathrm{NM}}$ is the average reciprocity in null model samples [garlaschelli2004patterns]. It measures the difference between observed and expected reciprocity by the maximum possible deviation, bounding $\rho_{\mathrm{NM}}$ between 1-1 and $1$. Positive values indicate more reciprocity than expected at random, negative values indicate less, and values near zero suggest consistency with the null model. This normalization allows a more faithful comparison and ranking of reciprocity across systems with different scales and density [garlaschelli2004patterns]. Recognizing its broad importance, recent works have extended reciprocity to hypergraphs, accounting for the complexity of having multiple nodes in both the source and target sets of hyperedges. Among the recent approaches for hypergraph reciprocity, one method decomposes hyperedges into pairwise links [pearcy2014hypergraph], losing information about group interactions. An alternative approach defines a more complex measure that diverges from the traditional binary definition of reciprocity at the level of single links [kim2023reciprocity]. While this approach can capture different nuances, it is computationally expensive and less straightforward to interpret, as it provides a continuous value instead of a simple yes-or-no answer to whether an interaction is reciprocated.
Figure 5: Higher-order reciprocity in real-world hypergraphs. a) Reciprocity score across datasets and reciprocity definitions. Each column corresponds to a distinct notion of higher-order reciprocity, thereby inducing a ranking of the datasets based on their scores. Datasets from the same domains share the same color. Arrows link the datasets across different definitions. b) Reciprocity score disaggregated by hyperedge size for each different notion of reciprocity. Trends should be interpreted relative to the null (zero line). To simplify the plots, we aggregate systems from the same domain.
@@
[…]
@@
[…]
We summarize our definitions of reciprocity for directed hypergraphs in Figure 4. More information about the algorithmic aspects of such measures is available in the Methods section.
After introducing these definitions, a natural first question is which systems exhibit the highest and lowest levels of reciprocity and how the ranking of systems based on reciprocity changes across different definitions. We address this in Figure 5a, which shows the normalized ratio of reciprocated hyperedges $\rho_{\mathrm{NM}}$ (reciprocity score) for each system across varying notions of reciprocity. The reciprocity score induces rankings of the systems, allowing us to observe which systems exhibit stronger tendencies toward mutual exchange of information. By definition, the score tends to increase for each system as we move from stricter definitions of reciprocity (exact) to more relaxed ones (weak). We observe that systems from the same domain tend to show similar levels of reciprocity across definitions, indicating that functional similarities within domains may drive comparable reciprocity patterns. E-mail datasets exhibit the highest levels of reciprocity, while Bitcoin datasets consistently show the lowest. Interestingly, while the ranking of systems remains largely stable with varying definitions, the relative distances between the datasets change. For instance, exact reciprocity mostly characterizes E-mail datasets, which are positioned far from the other datasets, clustering distinctly at the top of the scale. Strong reciprocity induces three clear clusters of datasets based on their scores: E-mail datasets rank the highest by a large margin, while Bitcoin datasets occupy the very low end. In the case of weak reciprocity, the datasets begin to separate along domain lines, spanning the entire spectrum of reciprocity scores. Notably, we observe a reduction in the distance between E-mail and datasets from metabolic and citation domains, suggesting a convergence in reciprocity levels as the definition becomes more relaxed. Overall, these patterns highlight how the choice of measure can influence the perceived level of reciprocity within different systems. By analyzing how the score evolves across definitions, we gain a more precise understanding of the extent of mutual exchange within each system, from the high reciprocity observed in E-mail datasets, where high mutual exchange is clear, to the lower reciprocity in Bitcoin datasets, where reciprocal connections are minimal across all definitions, and to the metabolic datasets, which emerge with high reciprocity under weaker definitions.
A related question is how the size of hyperedges influences the levels of reciprocity. We address this in Figure 5b, which reports the reciprocity score as a function of interaction size for the three notions, with deviations interpreted relative to the null baseline. For the E-mail datasets, $\rho_{\mathrm{NM}}$ displays consistent positive offset for strong and weak reciprocity across all admissible sizes, whereas exact reciprocity exhibits a pronounced signal only at size 2 and is negligible at larger sizes. This suggests that excess reciprocal structure relative to the null is present for most interaction sizes, while exact ”mirror” reversals remain largely a dyadic feature. In the Metabolic datasets, exact reciprocity remains near zero across sizes, with strong and weak deviations showing initial negative values for small sizes, followed by a substantial increase at higher sizes that is more pronounced for the weak definition. Thus, small-set interactions are under-reciprocated relative to the null, whereas large-set interactions become increasingly over-represented. In the Citation datasets, exact reciprocity is modestly positive at small sizes. Strong reciprocity displays a higher, stable excess across sizes, while weak reciprocity reaches its highest value for the smallest sets and diminishes with size. Citation reciprocity above the null is thus dominated by weak reciprocity in small interactions, with strong reciprocity contributing a more size-invariant offset. For the Bitcoin datasets, scores for all definitions cluster near zero, with a slight increase from exact to strong to weak, and minimal size dependence. This points to weak, size-stable excess reciprocal structure. Finally, the qna datasets track the Bitcoin pattern, with modest positive offsets that increase from exact to strong to weak definitions, and little systematic variation as size increases.
These findings suggest that weaker notions of reciprocity are valuable in providing insights into the overall reciprocity of systems with larger interactions, capturing a multifaceted view of how reciprocity operates at different strengths and interaction sizes.
@@
Motif analysis involves counting the frequency of patterns of interactions in connected subgraphs of a given number of nodes. This framework was first introduced by Milo et al. [milo2002network] to extract the fundamental functional units of complex systems [milo2004superfamilies]. Recently, motif analysis has been extended to undirected hypergraphs to capture patterns of interactions with arbitrary size [lotito2022higher]. Here, we extend such analysis to consider also the direction of the hyperedges involved in the patterns.
@@
First, it is interesting to study the combinatorics of the patterns of directed subhypergraphs. There is no simple closed-form formula for counting the number of possible directed higher-order motifs as a function of their order $n$, i.e., the number of nodes in the patterns. We can estimate the number of non-isomorphic connected directed hypergraphs in a way similar to [lotito2022higher]. Given a set of $n$ nodes, the number of possible directed hyperedges is $3^{n}-2\cdot\sum_{k=1}^{n}\binom{n}{k}-1=3^{n}-2\cdot 2^{n}+1$. This expression counts the ways to partition the $n$ nodes into three disjoint sets: source, target and empty set. We subtract the invalid combinations with empty source or target sets. Given $n$ nodes, we ensure connectivity by selecting a chain of n1n-1 hyperedges and including them in the hypergraph, leaving us with $3^{n}-2\cdot 2^{n}-n+2$ remaining possible hyperedges. For each remaining hyperedge, we decide whether to include it or not, resulting in $2^{3^{n}-2\cdot 2^{n}-n+2}$ total hypergraphs. Since we are interested in non-isomorphic hypergraphs, we divide this number by $n!$, the number of ways to label the vertices, providing the lower bound $\frac{2^{3^{n}-2\cdot 2^{n}-n+2}}{n!}$. If we ignore the constraints of non-isomorphism and connectivity, we count the number of possible labeled hypergraphs. Since each of the $3^{n}-2\cdot 2^{n}+1$ possible hyperedges can either be included or excluded, the total number of labeled hypergraphs is at most $2^{3^{n}-2\cdot 2^{n}+1}$. Figure 6 shows the upper and lower bounds on the growth of possible sub-hypergraph patterns as a function of the number of nodes (order) for both the undirected and directed cases. The estimated number of patterns grows super-exponentially, even in the undirected case. In the directed case, the growth is even faster due to the need to consider all possible subdivisions into source and target sets.
To perform motif analysis on real-world directed hypergraphs, we propose an exact algorithm to count the frequency of all connected sub-hypergraph patterns and associate each pattern with a z-score that quantifies its over- or under-representation relative to our proposed configuration model (see Methods).